OpenStudy (anonymous):

What is the center of the following ellipse? (x-5)^2 + (y+2)^2 ________ ________ = 1 25 36

OpenStudy (jdoe0001):

\(\bf \cfrac{(x-5)^2}{25}+\cfrac{(y+2)^2}{36}=1\implies \cfrac{(x-5)^2}{5^2}+\cfrac{(y-(-2))^2}{6^2}=1\\ \quad \\ \quad \\ \cfrac{(x-h)^2}{a^2}+\cfrac{(y-k)^2}{b^2}=1\qquad \textit{center is at }(h,k)\)

OpenStudy (anonymous):

Thank you!

OpenStudy (jdoe0001):

yw